OpenStudy (anonymous):

solve the following equations −8≤2x−4<2

OpenStudy (jdoe0001):

do you know how to simplify linear equations? like say solving for "x"?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

well. then that one is pretty much the same way in a linear equation you have 2 sides left side and right side in this one you just have 3 sides left < middle < right and anything you add or subtract or else to one, you have to do it to all three

OpenStudy (anonymous):

so can i add 4 to the 2

OpenStudy (jdoe0001):

\(\bf -8\le 2x-4\le 2\implies -8+4\le 2x\cancel{ -4+4 }\le 2+4\implies -4\le 2x\le 6 \\ \quad \\ \cfrac{\cancel{ -4 }}{\cancel{ 2 }}\le \cfrac{\cancel{ 2 }x}{\cancel{ 2 }}\le\cfrac{\cancel{ 6 }}{\cancel{ 2 }}\)

OpenStudy (anonymous):

so whats the answer

OpenStudy (jdoe0001):

well.. what would that end up as?

OpenStudy (jdoe0001):

say... what's left for the left-side the middle-side and the right-side

OpenStudy (anonymous):

-2 and 3

OpenStudy (jdoe0001):

\(\large \cfrac{\cancel{ -4 }}{\cancel{ 2 }}\le \cfrac{\cancel{ 2 }x}{\cancel{ 2 }}\le\cfrac{\cancel{ 6 }}{\cancel{ 2 }}\implies -2\le x\le 3\to \begin{cases} x\ge -2\\ x\le 3 \end{cases}\)